Shrewsbury Herbal Products, Ltd.

Shrewsbury Herbal Products, located in central England close to the Welsh border, is an old-line producer of herbal teas, seasonings, and medicines. Its products are marketed all over the United Kingdom and in many parts of continental Europe as well.

Shrewsbury Herbal generally invoices in British pound sterling when it sells to foreign customers in order to guard against adverse exchange rate changes. Nevertheless, it has just received an order from a large wholesaler in central France for £320,000 of its products, conditional upon delivery being made in three months’ time and the order invoiced in euros.

Shrewsbury’s controller, Elton Peters, is concerned with whether the pound will appreciate versus the euro over the next three months, thus eliminating all or most of the profit when the euro receivable is paid. He thinks this an unlikely possibility, but he decides to contact the firm’s banker for suggestions about hedging the exchange rate exposure.

Mr. Peters learns from the banker that the current spot exchange rate in €/£ is €1.4537; thus the invoice amount should be €465,184. Mr. Peters also learns that the three-month forward rates for the pound and the euro versus the U.S. dollar are $1.8990/£1.00 and $1.3154/€1.00, respectively. The banker offers to set up a forward hedge for selling the euro receivable for pound sterling based on the €/£ forward cross-exchange rate implicit in the forward rates against the dollar.

What would you do if you were Mr. Peters?
Answer #1

Since company is facing risk that that pound will appreciate versus euro it shall buy forward contract on pound or sell forward contract on euro

Contract price is euro 465184

Given the rates one should perform following

1)Sell euro- buy dollar

1 euro = 1.3154$

465184 =?

=465184*1.3154 = 611903.0336$

2) Sell dollar- buy pounds

1 pound = $1.8990

? = $611903.0336

=611903.0336/1.8990

=3,22,223.8197 pounds receivable

Thus by entering into hedge company will able to realize 3,22,223.8197 pounds

Thus its better to hedge as after hedging receivable amount is higher than originally invoiced amount